OpenStudy (micheal1234):

solve for t 32t = 122

OpenStudy (micheal1234):

divided by 32

OpenStudy (explorerhills):

what will you diveide on both sides so that you have just 't' on the RHS

OpenStudy (explorerhills):

yes good

OpenStudy (explorerhills):

t = 122/32 Just simplify

OpenStudy (anonymous):

t = 61/16 = 3.813

OpenStudy (micheal1234):

i got 3.48

OpenStudy (explorerhills):

how ? check the division part again?

OpenStudy (explorerhills):

and do you want t as a decimal or fraction?

OpenStudy (micheal1234):

decimal

OpenStudy (explorerhills):

ok.. 122/32 = 61/16 agreed?

OpenStudy (micheal1234):

yes

OpenStudy (explorerhills):

Now just divide..

OpenStudy (micheal1234):

is it 3.812

OpenStudy (explorerhills):

that's correct!

OpenStudy (explorerhills):

good work!